# Iteration Head: A Mechanistic Study of Chain-of-Thought

### v1

@@ 2 Controlled Setup: Learning Iterative Algorithms
We define iterative algorithms, or iterative schemes, as follows: an iterative algorithm is the combination of an input sequence, denoted as Sequence, and made of $L$ elements $(x_{t})_{t\in[L]}$ (with $[L]=\{1,\cdots,L\}$), and an internal state, denoted as $s$, initialized to some default value $s_{0}=\texttt{Init}$, and updated as the sequence is processed according to some rule $s_{t}=F(s_{t-1},x_{t})$ for some function $F$. Pseudo-code illustrating the concept is provided by Algorithm 1. By extension, we informally call iterative task a task which is naturally solved by outputting the end product of some iterative algorithm applied to some input sequence. As an example, consider the parity problem, i.e. the problem of computing the parity of the sum of a sequence of 00s and $1$s: it can be easily framed as an iterative task. Using the notations of Algorithm 1, let the initial state Init be equal to 00, and let $F(s,x)$ be equal to 00 if $s$ is equal to $x$, and $1$ otherwise. Then the final $s_{L}$ gives the parity of the sum. Although this task could also be solved in a non-iterative fashion, the iterative solution can be seen as simpler and more parsimonious.
@@
Given a certain number of transformer blocks, a transformer can only apply a corresponding number of cross-operations to predict the next token. This limits its ability to learn even relatively simple iterative tasks. E.g., consider the task where the input sequence is $(x_{1},\ldots,x_{L})$, possibly restricted to $x_{i}\in[a,b]$ for some $a<b$, and the desired output is the product $\prod x_{i}$. The product is multilinear in the entries of the input sequence $(x_{1},\ldots,x_{L})$. If we model the output of an attention layer as sums of monomials of degree at most three in its input variables (due to key-query-value interaction), this makes learning the task quite hard for a transformer, and bounds the maximum length of the sequences that a transformer with a given number of blocks can correctly process [see 51, and related literature for formal discussions on the matter that capture this log-depth dependency].
However, when transformers are allowed to generate many tokens before providing an answer, which implicitly lifts the constraint on the number of operations performed by the transformer, the picture changes [48, 39, 35] [see also 14, 21]. In particular, Figure 1, explained in the next subsection, illustrates how a two-layer transformer can implement what we named an “iteration head”. This potentially enables it to learn any iterative algorithm, assuming that its second layer MLP is big enough to implement any successor function $F:(x_{t},s_{t-1})\mapsto s_{t}$.
@@ 3.1 Theoretical Circuit
Iteration heads are an efficient, flexible and parsimonious way to implement iterative algorithms; as such, we expect them to naturally emerge during training. Nonetheless, transformers have flexible architectures that can perform similar operations in different ways. Hence we also expect to see some variations with respect to the schematic architecture described above (see Figures 4 and 9), in particular when the embedding dimension becomes too small for the information superposition from the previous paragraph to be correctly implemented.
@@ 3.2 Learning an Iteration Head
Unless otherwise stated, our experimental setup is as follows. Data was generated for the binary-copy, parity, and polynomial iteration problem with $P(X,Y)=XY+1$ in ${\mathbb{F}}_{11}$. For each length $L$ from $L_{\min}=1$ to $L_{\max}=32$, we generated $n=1024$ input sequences of length $L$ (corresponding to a total sequence length of $32L+$ 3) uniformly at random for both training and testing sets, creating datasets of $N=16,384=16\times 1024$ sequences in total. [^2] Note that, even when $x_{t}\in{\mathbb{F}}_{2}$, the expected percentage of collision between the training and testing set decreases exponentially fast with $L$, ensuring minimal contamination between training and testing. We utilized auto-regressive transformers [10] with two layers and one attention head per layer. The embedding dimension was set to $d=128$, with learned absolute positional encoding added to the learned token embedding. The weights were optimized over 1000 epochs with Adam [29], a batch size of 256, and a fixed learning rate set to $\gamma=3\cdot 10^{-4}$, with default PyTorch parameters otherwise [46]. Our experiments consumed 12k V100-hours.
@@ 4.2 The Role of Inductive Biases
Figure 6: Same figure as Figure 3, except that we considered the parity problem, and 5000 training epochs.
@@ 5 Conclusion
Societal Impact.
Mechanistic interpretability is focused on understanding the key mechanisms at play in deep learning systems by tracing them down to the weights. It is often associated with AI safety, hoping that a deeper understanding of these systems can help us steer them to be more “aligned” with “human values”, and prevent AI dystopia scenarios. On the other hand, it could also prove useful in training more powerful models, which is associated with significant societal issues linked to the rise of advanced AI systems. These issues are too broad to be discussed in this paragraph.
@@ Appendix A Additional Figures & Findings
When we started this project, we were expecting to find some grokking structure emerged from the need to perform position subtraction. In particular, as mentionned in the main text, we were expecting this mechanism to appear as the position embedding dimension was small. When learning on the parity problem, we found that the network was not implementing the iteration head when the position embedding dimension was really small. This can notably be seen on Figures 11 and 10. We notably observed that freezing the position embedding does not change much the picture, which can be seen as a result of overparameterization. Similar type of observation were observed when learning with the polynomial iteration problem, as reported on Figure 12.

### v2

@@ 2 Controlled Setup: Learning Iterative Algorithms
We define iterative algorithms, or iterative schemes, as follows: an iterative algorithm is the combination of an input sequence, denoted as Sequence, and made of $L$ elements $(x_{t})_{t\in[L]}$ (with $[L]=\{1,\cdots,L\}$), and an internal state, denoted as $s$, initialized to some default value $s_{0}=\texttt{Init}$, and updated as the sequence is processed according to some rule $s_{t}=F(s_{t-1},x_{t})$ for some function $F$. Pseudo-code illustrating the concept is provided by Algorithm 1, see also Figure 4. By extension, we informally call iterative task a task which is naturally solved by outputting the end product of some iterative algorithm applied to some input sequence. As an example, consider the parity problem, i.e. the problem of computing the parity of the sum of a sequence of 00s and $1$s: it can be easily framed as an iterative task. Using the notations of Algorithm 1, let the initial state Init be equal to 00, and let $F(s,x)$ be equal to 00 if $s$ is equal to $x$, and $1$ otherwise. Then the final $s_{L}$ gives the parity of the sum. Although this task could also be solved in a non-iterative fashion, the iterative solution can be seen as simpler and more parsimonious.
$x_{t-3}$$x_{t-2}$$x_{t-1}$$x_{t}$$s_{t-1}$$s_{t}$(Tokens)(Internal States)
Figure 1: Arguably, reasoning involves updating an internal state (red) as new information is processed (green). The diagram above, where each element represents a piece of information, is an abstract depiction of this idea. This observation motivates our use of iterative tasks as a proxy for more general reasoning processes. At first glance, a limitation of transformers is their lack of an internal state, which makes it challenging to implement this diagram [32].
$x_{t-3}$$x_{t-2}$$x_{t-1}$$x_{t}$$s_{t-1}$$x_{t}$$s_{t}$(Tokens)(Internal Layer)(Additional Layer)
Figure 2: A single transformer layer cannot implement the diagram from 4, as it cannot access its previous outputs. This limitation can be bypassed by stacking transformer layers, as illustrated here. The red arrow indicates a residual connection. This naive method requires as many layers as there are reasoning hops.
$x_{t-3}$$x_{t-2}$$x_{t-1}$$x_{t}$$s_{t-1}$$s_{t}$(Tokens)(Internal Layer)
Figure 3: Alternatively, a transformer could compute each state $s_{t}$ from scratch. This implementation does not require additional layers, but it is not parsimonious, which could lead to computational inefficiencies. This explains the difficulty for a transformer to output the final answer of a chain of reasoning within a token (i.e., with next-token prediction, and without chain-of-thought)
[…]
Figure 4: Chain-of-thought addresses this issue by explicitly representing the reasoning process in token space. The auto-regressive nature of LLMs (blue arrow) allows for the implementation of iterative algorithms, as long as the states are encoded in token space. A concrete implementation of such a mechanism, which we call iteration head, is described in Section 3. In practical applications of LLMs, one could imagine earlier layers summarizing the $t$-th input sentence (or some other coherent semantic information of varying token length) into $x_{t}$, as well as summarizing the generated CoT sentences into some $s_{t-1}$, with later layers translating the state $s_{t}$ into readable text.
@@
Given a certain number of transformer blocks, a transformer can only apply a corresponding number of cross-operations to predict the next token. This limits its ability to learn even relatively simple iterative tasks, see Figures 4 and 4. E.g., consider the task where the input sequence is $(x_{1},\ldots,x_{L})$, possibly restricted to $x_{i}\in[a,b]$ for some $a<b$, and the desired output is the product $\prod x_{i}$. The product is multilinear in the entries of the input sequence $(x_{1},\ldots,x_{L})$. If we model the output of an attention layer as sums of monomials of degree at most three in its input variables (due to key-query-value interaction), this makes learning the task quite hard for a transformer, and bounds the maximum length of the sequences that a transformer with a given number of blocks can correctly process [see 51, and related literature for formal discussions on the matter that capture this log-depth dependency].
However, when transformers are allowed to generate many tokens before providing an answer, which implicitly lifts the constraint on the number of operations performed by the transformer (see Figure 4), the picture changes [48, 39, 35] [see also 14, 21]. In particular, Figure 5, explained in the next subsection, illustrates how a two-layer transformer can implement what we named an “iteration head”. This potentially enables it to learn any iterative algorithm, assuming that its second layer MLP is big enough to implement any successor function $F:(x_{t},s_{t-1})\mapsto s_{t}$.
@@ 3.1 Theoretical Circuit
Iteration heads are an efficient, flexible and parsimonious way to implement iterative algorithms; as such, we expect them to naturally emerge during training. Nonetheless, transformers have flexible architectures that can perform similar operations in different ways. Hence we also expect to see some variations with respect to the schematic architecture described above (see Figures 8 and 13), in particular when the embedding dimension becomes too small for the information superposition from the previous paragraph to be correctly implemented.
@@ 3.2 Learning an Iteration Head
Unless otherwise stated, our experimental setup is as follows. Data was generated for the binary-copy, parity, and polynomial iteration problem with $P(X,Y)=XY+1$ in ${\mathbb{F}}_{11}$. For each length $L$ from $L_{\min}=1$ to $L_{\max}=32$, we generated $n=1024$ input sequences of length $L$ (corresponding to a total sequence length of $32L+$ 3) uniformly at random for both training and testing sets, creating datasets of $N=16,384=16\times 1024$ sequences in total. [^2] Note that, even when $x_{t}\in{\mathbb{F}}_{2}$, the expected percentage of collision between the training and testing set decreases exponentially fast with $L$, ensuring minimal contamination between training and testing. We utilized auto-regressive transformers [10] with two layers and one attention head per layer. The embedding dimension was set to $d=128$, with learned absolute positional encoding added to the learned token embedding. The weights were optimized over 1000 epochs with Adam [29], a batch size of 256, and a fixed learning rate set to $\gamma=3\cdot 10^{-4}$, with default PyTorch parameters otherwise [46]. Our source code is available at https://github.com/facebookresearch/pal. Our experiments consumed 12k V100-hours.
@@ 3.3 Ablation Studies
In addition to visualizing the attention map, we validated the learning of iteration heads through attention patching, i.e., intervening to “patch” certain attention maps. Specifically, we observed that patching the ideal attention maps (i.e., zeroing out other routes) does not disrupt perfect accuracy. In contrast, zeroing out the focus on the EoI by the first attention head, or on $p_{t}$ by the second, reduced performance to near random.
@@ 4.2 The Role of Inductive Biases
Figure 10: Same as Figure 7, except that we considered the parity problem, and 5000 training epochs.
@@ Appendix A Additional Figures & Findings
When we started this project, we were expecting to find some grokking structure emerged from the need to perform position subtraction. In particular, as mentionned in the main text, we were expecting this mechanism to appear as the position embedding dimension was small. When learning on the parity problem, we found that the network was not implementing the iteration head when the position embedding dimension was really small. This can notably be seen on Figures 15 and 14. We notably observed that freezing the position embedding does not change much the picture, which can be seen as a result of overparameterization. Similar type of observation were observed when learning with the polynomial iteration problem, as reported on Figure 16.
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We composed the abstract and introduction after establishing the main results.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations of our synthetic data study should be clearly apparent to the reader.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The results were stated after being demonstrated. Some results that were obvious to us were not demonstrated (assuming one does not have to go back to Peano axiomatic to write additions).
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The codebase contains all the necessary details, and random seeds were set to ensure full reproducibility.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The code is freely available.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The codebase is available with all the training and testing details. Some details (e.g., pre-norm, embedding dropout rate) were omitted from the main paper to prevent overwhelming the reader with technicalities that do not form the core of the paper.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Please note that in some 2D plots, “error bars” can be inferred from the texture of the contour lines, which provide a clear sense of the randomness in the process.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We briefly comment on the resources used. The code is open source and the experiments are relatively light. An interested reader could easily calculate a more precise estimate of the energy cost to reproduce the results.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Adherence to guidelines was of significant importance to us.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: This work is focused on enhancing our understanding of machine learning systems. While this endeavor comes with clear issues, they are too broad for us to provide any meaningful comments within this context.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper was deemed too theoretical to necessitate significant consideration regarding safety concerns.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We didn’t cite everything (for instance, we didn’t mention Python or CUDA), but we did cite PyTorch, the transformer paper, as well as GPT-3.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We have documented our synthetic datasets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We have not conducted any experiments involving human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: See above.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.